# Empirical Bayes Estimation in Heterogeneous Coefficient Panel Models

### v1

@@ 5 Regret Consistency and $\ell_{p}$ Convergence
Let $(\tau_{i})_{i=1}^{N}=(\tau(\theta_{i}))_{i=1}^{N}$ denote the unit-level latent parameters of interest, each defined as a function of $\theta_{i}$. [^7] We focus on the case of a scalar parameter here. However, the discussion and results in this section readily extend to the vector-valued case. The estimation problem consists of constructing a decision rule $(Y_{i},X_{i})_{i=1}^{N}\mapsto(\hat{\tau}_{i})_{i=1}^{N}$. Under the compound decision framework, the econometrician seeks to minimize the compound squared-error loss, defined as
$L(\hat{\tau},\tau):=\frac{1}{N}\sum_{i=1}^{N}(\hat{\tau}_{i}-\tau_{i})^{2}.$
The compound risk $R(\hat{\tau},\tau):=\mathbb{E}_{G_{*}}[L(\hat{\tau},\tau)]$ is obtained by taking the expectation of the compound loss jointly over $(Y_{i},X_{i})_{i=1}^{N}$ and $(\theta_{i})_{i=1}^{N}$, accounting for the dependence of the estimator $(\hat{\tau}_{i})_{i=1}^{N}$ on the data. It is well-known that the squared-error risk is minimized by the oracle posterior mean:
@@
where the second equality follows from the independence across units. In contrast to the separate decision rule $\hat{\tau}^{\mathrm{MLE}}_{i}=\tau(\hat{\theta}^{\operatorname{MLE}}_{i})$, efficiency gains are achieved by pooling information across units through $G_{*}$. G-modeling approximates the oracle (14) by substituting $\hat{G}$ for $G_{*}$, yielding the estimator
This section establishes asymptotic guarantees for the EB method in terms of regret, defined as:
@@
since $\mathbb{E}_{G_{*}}[(\hat{\tau}^{\mathrm{EB}}_{i}-{\tau}^{*}_{i})({\tau}^{*}_{i}-\tau_{i})]=0$ by the properties of conditional expectation. The condition $\lim_{N\to\infty}[R(\hat{\tau}^{\mathrm{EB}},\tau)-R({\tau}^{*},\tau)]=0$ is referred to as regret consistency (see, e.g., Adusumilli, Gu, and Tao (2025)) or asymptotic optimality (see, e.g., Zhang (1997)). We mainly refer to this condition as regret consistency.
@@
The EB one-period-ahead prediction $\hat{\tau}^{\mathrm{EB}}_{i}$ is then obtained by substituting the NPMLE for $G_{*}$ in (15). Evaluating the EB method in this setting requires controlling the compound error arising from the interaction between the estimation discrepancy $\hat{\theta}^{\mathrm{EB}}_{i}-{\theta}^{*}_{i}$ and these observed variables. Consequently, the analysis necessitates controlling the $\ell_{p}$ distance (where $p\geq 2$, with $p=4$ being a particularly important case) between the EB and oracle estimates. We now establish regret consistency under the high-level assumptions stated below.
@@
Theorem 3 provides the theoretical justification for using the NPMLE-based EB rule in large samples. The primary challenge in establishing this result is that the weak convergence of the priors, $d(\hat{G},G_{*})\overset{\operatorname{a.s.}}{\longrightarrow}0$, implies the convergence of integrals only for bounded continuous functions. However, $\tau(\theta)$ and the squared-error loss are generally unbounded functions of the parameters. Assumption 4(ii) is critical because uniform integrability of the estimated prior moments ensures uniform integrability of the posterior moments across units. This allows us to upgrade the weak convergence of $\hat{G}$ to the convergence of the posterior expectations (the EB estimators) in $\ell_{p}$-norm. Notably, the second result of the theorem, that is, convergence in $\ell_{p}$-norm, is stronger than simple regret consistency ($p=2$) and is particularly useful when the latent parameters enter the decision problem nonlinearly or interact with unbounded covariates.
@@ 5.1 Regret Consistency in the HIVDX Model
When $\tau_{i}=\theta_{i}$, the EB estimator of $\theta_{i}$ is given by
$\hat{\theta}^{\mathrm{EB}}_{i}:=\frac{\int_{\Theta}\theta\ell(Y_{i}\mid X_{i},\theta)d\hat{G}(\theta)}{f_{\hat{G}}(Y_{i},X_{i})},$
which can be partitioned as $\hat{\theta}^{\mathrm{EB}}_{i}=(\hat{\beta}^{\mathrm{EB}}_{i},\hat{\delta}^{\mathrm{EB}}_{i})$. The regret consistency of $\hat{\delta}^{\mathrm{EB}}_{i}$ follows immediately from Theorem 3 and Assumption 4, as $\delta_{i}$ lies in a compact set. For $\hat{\beta}^{\mathrm{EB}}_{i}$, however, Assumption 4(ii) requires $\|\beta_{i}\|^{p}$ to be uniformly integrable with respect to the sequence of estimated priors. This condition is guaranteed by a moment condition on $(Y_{i},X_{i})$.
Once $(\theta_{i})_{i=1}^{N}$ have been estimated, we turn to predicting future outcomes. We focus on the one-period-ahead EB prediction $\hat{Y}^{\mathrm{EB}}_{iT+1}$ of $Y_{iT+1}$ (recall the oracle estimator given in (15)). Let $\bar{X}_{2,i}=T^{-1}\sum_{t=1}^{T}X_{2,it}$ denote the time average of $X_{2,it}$ for unit $i$.
@@
Let Assumption 1 hold. Assume that $|\bar{X}_{2,i}|\leq M$ and $\sum_{t=1}^{T}(X_{2,it}-\bar{X}_{2,i})^{2}\geq c$ for some constants $M,c>0$ and for all $i\geq 1$. Then, we have the following regret consistency results:
@@ 6 Computational Algorithms
A variety of computational methods have been proposed for implementing the NPMLE. Although the optimization problem is convex, computation is challenging because the optimization variable $G$ is an infinite-dimensional object, rendering direct implementation intractable. A common remedy is to approximate $G$ by a discrete distribution supported on $m$ atoms (with $m\in\mathbb{N}$), thereby reducing the problem to fitting an $m$-component mixture model. [^8] This approach is justified by the well-known result that a discrete NPMLE with at most $N_{0}$ support points exists (possibly among multiple maximizers), where $N_{0}$ is the number of distinct data points (Lindsay, 1983). Furthermore, Shen and Wu (2025) show that the required number of components $m$ grows at the rate $O(\log N)$ in the one-dimensional case under subgaussian priors, and conjecture that in higher dimensions $m$ increases on the order of $O((\log N)^{C})$, where $C$ depends on the dimension of $\theta_{i}$. For the specific case of one-dimensional Gaussian location mixtures, Polyanskiy and Sellke (2025) establish computational guarantees, including an algorithm that computes the exact support size.
For example, Jiang and Zhang (2009) introduce a fixed-point EM algorithm that mirrors the standard EM routine and establish that it attains a global optimum up to an explicit error bound. In practice, however, EM with a large fixed grid often converges prohibitively slowly, making it unsuitable for our HC model. Koenker and Mizera (2014) develop a convex-optimization method based on an interior-point algorithm to fit weights on a fixed grid, which Gu and Koenker (2017) apply to a dynamic panel model with two-dimensional heterogeneity. While generally reliable, this approach requires a prespecified grid whose size grows exponentially with the dimension of the support. More recently, Kim, Carbonetto, Stephens, and Anitescu (2020), Zhang, Cui, Sen, and Toh (2024), and Wang, Ibrahim, and Mazumder (2025) propose alternative methods based on general-purpose convex programming to improve scalability. However, high-dimensional computational challenges persist. For instance, Zhang, Cui, Sen, and Toh (2024) suggest using observed data points as the grid when the dimension of multivariate mixture models is three or higher.
A feature shared by all the methods above is that the location of the support points is fixed; optimization is performed solely with respect to the weights. In our HIVDX model ($d_{\theta}=4$), this imposes a severe trade-off: a grid dense enough to minimize approximation error becomes computationally intractable due to the curse of dimensionality, while a sparse grid risks significant discretization bias. Moreover, unlike in simple mixture models, there are no natural candidates (such as observed data points) to serve as a baseline grid in our setting. Consequently, these recent proposals are not well-suited for our purposes. Motivated by these challenges for multiple sources of heterogeneity, we consider a discrete-time Wasserstein-Fisher-Rao (WFR) gradient flow algorithm on the space of probability measures over $\Theta$. Crucially, the WFR algorithm adaptively relocates support points; as a result, the number of support atoms $m$ remains moderate and need not exceed the sample size $N$. We now explain our procedure in detail.
@@
Let $\boldsymbol{\theta}=(\theta^{j})_{j=1}^{m}$ denote a grid of support points for $G$, and let $\boldsymbol{\omega}=(\omega^{j})_{j=1}^{m}$ be the associated weights satisfying $\omega^{j}\geq 0$ for all $j$ and $\sum_{j=1}^{m}\omega^{j}=1$. We use the superscript $j$ to distinguish the grid $(\theta^{j})_{j=1}^{m}\in\Theta^{m}$ from the individual-specific parameters $(\theta_{i})_{i=1}^{N}\in\Theta^{N}$. Let $G^{\boldsymbol{\theta},\boldsymbol{\omega}}$ denote the discrete distribution induced by $(\boldsymbol{\theta},\boldsymbol{\omega})$, i.e.,
@@
$G^{\boldsymbol{\theta},\boldsymbol{\omega}}(\Theta^{\prime})=\sum_{j=1}^{m}\omega^{j}\,{\mathbf{1}\large[{\theta^{j}\in\Theta^{\prime}}\large]}\qquad\text{for all subsets }\Theta^{\prime}\subseteq\Theta.$
The NPMLE of the HC model can then be expressed as the solution to
@@
By the Karush-Kuhn-Tucker theorem, for some $\lambda\geq 0$, this problem is equivalent to
$\operatornamewithlimits{arg\hskip 1.00006ptmax}_{\boldsymbol{\theta}\in\Theta^{m},\,\boldsymbol{\omega}\geq 0}\frac{1}{N}\sum_{i=1}^{N}\log\left(\sum_{j=1}^{m}\ell(Y_{i}\,|\,X_{i},\theta^{j})\omega^{j}\right)-\lambda\sum_{j=1}^{m}\omega^{j},$ (16)
where the value of $\lambda=1$ can be determined by investigating the first-order conditions for each weight $\omega^{j}$.
Even with discretization, finding a path to the global maximizer remains challenging because the space of probability measures denoted by $\mathcal{P}(\Theta)$ has a complex geometry, and different metrics can generate markedly different gradient-flow trajectories. [^9] Gradient flow generalizes gradient descent to metric spaces beyond the standard Euclidean setting. In our case, the objective is the likelihood function, and the decision variable $G$ lies in the space of probability measures. For example, the commonly used Fisher-Rao (FR) metric and the quadratic Wasserstein (W) distance induce distinct notions of gradient flow on $\mathcal{P}(\Theta)$. The FR gradient flow updates $G_{t}^{\mathrm{FR}}$ according to
@@
where the reweighting strategy $\alpha_{t}:\Theta\to\mathbb{R}$ is chosen to maximize $F_{N}(G_{t+\Delta t}^{\mathrm{FR}})-F_{N}(G_{t}^{\mathrm{FR}})$ subject to the normalization $\mathbb{E}_{G_{t}^{\mathrm{FR}}}[(\alpha_{t}-\mathbb{E}_{G_{t}^{\mathrm{FR}}}[\alpha_{t}])^{2}]=1$. The limiting path $(G_{t}^{\mathrm{FR}})_{t\geq 0}$ as $\Delta t\to 0$ defines the FR gradient flow. In contrast, the W gradient flow evolves according to the continuity (transport) equation
@@
where the velocity field $V_{t}:\Theta\to\mathbb{R}^{{d_{\theta}}}$ describes the movement of mass and is chosen to maximize the gain in $F_{N}$ under the normalization $\mathbb{E}_{G_{t}^{\mathrm{W}}}[\|V_{t}\|^{2}]=1$. The resulting trajectory $(G_{t}^{\mathrm{W}})_{t\geq 0}$ is referred to as the W gradient flow. Because they are derived from different metrics, the FR and W flows yield different steady-state conditions. Notably, the first-order condition of the FR flow implies that of the W flow. The FR flow is guaranteed to reach a global maximum when initialized with full support, as established by Yan, Wang, and Rigollet (2024) for the Gaussian location mixture case, whereas the W flow may stall at a local maximum. Intuitively, the FR flow emphasizes reweighting existing support points, while the W flow emphasizes transporting mass across the space, leading to different convergence behavior.
The Wasserstein-Fisher-Rao (WFR) gradient flow combines both reweighting and transport dynamics when updating $G_{t}^{\mathrm{WFR}}$:
@@
As a result, the WFR gradient flow inherits the global convergence guarantees of the FR flow, while the W component allows it to adapt the support through mass transport. In the context of Gaussian location mixtures, Yan, Wang, and Rigollet (2024) provide theoretical guarantees for a discrete-time WFR gradient-descent algorithm and empirically demonstrates its superiority over both the EM algorithm and gradient-descent methods based solely on the FR or W flow. We extend their approach to a more general HC model that includes the HIVD and HIVDX models as special cases. Intuitively, WFR balances exploration through transport with refinement through reweighting, which explains its superior performance in both theory and practice. Further details are presented in Section 6.1, with a comparison to the EM algorithm in Section 6.2.

### v2

@@ 5 Regret Consistency and $\ell_{p}$ Convergence
Let $(\tau_{i})_{i=1}^{N}=(\tau(\theta_{i}))_{i=1}^{N}$ denote the unit-level latent parameters of interest, each defined as a function of $\theta_{i}$. [^7] We focus on the case of a scalar parameter here. However, the discussion and results in this section readily extend to the vector-valued case. The estimation problem consists of constructing a compound decision rule $(Y_{i},X_{i})_{i=1}^{N}\mapsto(\hat{\tau}_{i})_{i=1}^{N}$. G-modeling yields the estimator:
This section establishes asymptotic guarantees for $\hat{\tau}^{\mathrm{EB}}_{i}$ in terms of regret $R(\hat{\tau}^{\mathrm{EB}},\tau)-R({\tau}^{*},\tau)$, where $R(\hat{\tau},\tau):=\mathbb{E}_{G_{*}}[L(\hat{\tau},\tau)]$ is compound risk. Specifically, we consider quadratic loss $L(\hat{\tau},\tau):=\frac{1}{N}\sum_{i=1}^{N}(\hat{\tau}_{i}-\tau_{i})^{2}$, and hence the oracle posterior mean is defined by:
@@
where the second equality follows from the independence across units. Write
@@
since $\mathbb{E}_{G_{*}}[(\hat{\tau}^{\mathrm{EB}}_{i}-{\tau}^{*}_{i})({\tau}^{*}_{i}-\tau_{i})]=0$ by the properties of conditional expectation. The condition $\lim_{N\to\infty}[R(\hat{\tau}^{\mathrm{EB}},\tau)-R({\tau}^{*},\tau)]=0$ is referred to as regret consistency (see, e.g., Adusumilli, Gu, and Tao (2025)) or asymptotic optimality (see, e.g., Zhang (1997)).
@@
The EB one-period-ahead prediction $\hat{\tau}^{\mathrm{EB}}_{i}$ is then obtained by substituting the NPMLE for $G_{*}$ in (14). Evaluating the EB method in this setting requires controlling the compound error arising from the interaction between the estimation discrepancy $\hat{\theta}^{\mathrm{EB}}_{i}-{\theta}^{*}_{i}$ and these observed variables. Consequently, the analysis necessitates controlling the $\ell_{p}$ distance (where $p\geq 2$, with $p=4$ being a particularly important case) between the EB and oracle estimates.
@@
Theorem 3 provides the theoretical justification for using $(\hat{\tau}^{\mathrm{EB}}_{i})_{i=1}^{N}$ in large samples. The primary challenge in establishing this result is that the weak convergence of the priors, $d(\hat{G},G_{*})\overset{\operatorname{a.s.}}{\longrightarrow}0$, implies the convergence of integrals only for bounded continuous functions. However, $\tau(\theta)$ and the squared-error loss are generally unbounded functions of the parameters. Assumption 4(ii) is critical because uniform integrability of the estimated prior moments ensures uniform integrability of the posterior moments across units. This allows us to upgrade the weak convergence of $\hat{G}$ to the convergence of the posterior expectations (the EB estimators) in $\ell_{p}$-norm. Notably, the second result of the theorem, that is, convergence in $\ell_{p}$-norm, is stronger than simple regret consistency ($p=2$) and is particularly useful when the latent parameters enter the decision problem nonlinearly or interact with unbounded covariates.
@@ 5.1 Regret Consistency in the HIVDX Model
When $\tau_{i}=\theta_{i}$ and $\theta_{i}=(\beta_{i}^{\prime},\delta_{i}^{\prime})^{\prime}$ in the HIVDX model, the EB estimator of $\theta_{i}$ can be partitioned as $\hat{\theta}^{\mathrm{EB}}_{i}=(\hat{\beta}^{\mathrm{EB}}_{i},\hat{\delta}^{\mathrm{EB}}_{i})$. Regret consistency of $\hat{\delta}^{\mathrm{EB}}_{i}$ follows immediately from Theorem 3 and Assumption 4, as $\delta_{i}$ lies in a compact set. For $\hat{\beta}^{\mathrm{EB}}_{i}$, however, Assumption 4(ii) requires $\|\beta_{i}\|^{p}$ to be uniformly integrable with respect to the sequence of estimated priors. This condition is guaranteed by a moment condition on $(Y_{i},X_{i})$.
Once $(\theta_{i})_{i=1}^{N}$ have been estimated, we turn to the one-period-ahead EB prediction $\hat{Y}^{\mathrm{EB}}_{iT+1}$ of $Y_{iT+1}$. Given the oracle estimator in (14), the prediction error involves interaction of the errors in estimating the random coefficients and $(Y_{i},X_{i})$.
Let $\bar{X}_{2,i}=T^{-1}\sum_{t=1}^{T}X_{2,it}$ denote the time average of $X_{2,it}$ for unit $i$.
@@
Let Assumption 1 hold. Assume that $|\bar{X}_{2,i}|\leq M$ and $\sum_{t=1}^{T}(X_{2,it}-\bar{X}_{2,i})^{2}\geq c$ for some constants $M,c>0$ and for all $i\geq 1$.
@@ 6 Computational Algorithms
A variety of computational methods have been proposed for implementing the NPMLE. Although the optimization problem is convex, computation is challenging because the optimization variable $G$ is an infinite-dimensional object, rendering direct implementation intractable. A common remedy is to approximate $G$ by a discrete distribution supported on $m$ atoms (with $m\in\mathbb{N}$), thereby reducing the problem to fitting an $m$-component mixture model. This approach is justified by the well-known result that a discrete NPMLE with at most $N_{0}$ support points exists (possibly among multiple maximizers), where $N_{0}$ is the number of distinct data points (Lindsay, 1983). [^8] Shen and Wu (2025) show that the required number of components $m$ grows at the rate $O(\log N)$ in the one-dimensional case under subgaussian priors, and conjecture that in higher dimensions $m$ increases on the order of $O((\log N)^{C})$, where $C$ depends on the dimension of $\theta_{i}$. For one-dimensional Gaussian location mixtures, Polyanskiy and Sellke (2025) establish computational guarantees, including an algorithm that computes the exact support size.
@@
Specifically, let $\boldsymbol{\theta}=(\theta^{j})_{j=1}^{m}$ denote a grid of support points for $G$, and let $\boldsymbol{\omega}=(\omega^{j})_{j=1}^{m}$ be the associated weights satisfying $\omega^{j}\geq 0$ for all $j$ and $\sum_{j=1}^{m}\omega^{j}=1$. Let $G^{\boldsymbol{\theta},\boldsymbol{\omega}}$ denote the discrete distribution induced by $(\boldsymbol{\theta},\boldsymbol{\omega})$, i.e.,
@@
$G^{\boldsymbol{\theta},\boldsymbol{\omega}}(\Theta^{\prime})=\sum_{j=1}^{m}\omega^{j}\,{\mathbf{1}\large[{\theta^{j}\in\Theta^{\prime}}\large]}\qquad\text{for all subsets }\Theta^{\prime}\subseteq\Theta,$
where we use the superscript $j$ to denote the grid $(\theta^{j})_{j=1}^{m}\in\Theta^{m}$, as distinct from the individual-specific parameters $(\theta_{i})_{i=1}^{N}\in\Theta^{N}$. The NPMLE of the HC model is given by
@@
By the Karush-Kuhn-Tucker theorem, this problem is equivalent to maximizing the Lagrangian
$\mathcal{L}(\boldsymbol{\theta},\boldsymbol{\omega})=\frac{1}{N}\sum_{i=1}^{N}\log\left(\sum_{j=1}^{m}\ell(Y_{i}\,|\,X_{i},\theta^{j})\omega^{j}\right)-\lambda\left(\sum_{j=1}^{m}\omega^{j}-1\right).$ (15)
The first-order conditions with respect to $\boldsymbol{\omega}$ imply that the optimal Lagrange multiplier is fixed at $\lambda=1$.
Most EB estimators are solutions to the discretized problem described above. For example, Jiang and Zhang (2009) introduce a fixed-point EM algorithm that mirrors the standard EM routine and establish that it attains a global optimum up to an explicit error bound. Gu and Koenker (2017) estimate a dynamic panel model with two-dimensional heterogeneity using the interior-point algorithm of Koenker and Mizera (2014) to fit weights on a fixed grid. More recently, Kim, Carbonetto, Stephens, and Anitescu (2020), Zhang, Cui, Sen, and Toh (2024), and Wang, Ibrahim, and Mazumder (2025) propose alternative methods based on general-purpose convex programming to improve scalability.
A feature shared by all these methods is that the location of the support points is fixed, so optimization is performed solely with respect to the weights $\boldsymbol{\omega}$. Unlike in simple mixture models, determining where to position the grid points is not obvious when the dimension of $\theta$ exceeds two. Zhang, Cui, Sen, and Toh (2024) suggest using observed data points as the grid when the dimension is three or higher. The choice of grid size is also delicate: a grid dense enough to minimize approximation error becomes computationally intractable due to the curse of dimensionality, while a sparse grid risks significant discretization bias.
To overcome the limitations of fixed grids, we consider an algorithm that allows the support points to move adaptively. This necessitates a shift from optimizing weights on a static grid to optimizing the probability distribution $G$ itself within the space of probability measures $\mathcal{P}(\Theta)$. In this infinite-dimensional setting, the standard gradient descent used in Euclidean spaces is replaced by gradient flow, where the specific trajectory depends critically on the geometric structure imposed on $\mathcal{P}(\Theta)$. Because $\mathcal{P}(\Theta)$ can have a complex geometry, different metrics can generate markedly different gradient-flow trajectories.
Consider the gradient flow induced by two commonly used metrics: the Fisher-Rao (FR) and the Wasserstein (W). The FR gradient flow updates $G_{t}^{\mathrm{FR}}$ according to
@@
where the reweighting strategy $\alpha_{t}:\Theta\to\mathbb{R}$ is chosen to maximize the likelihood gain $F_{N}(G_{t+\Delta t}^{\mathrm{FR}})-F_{N}(G_{t}^{\mathrm{FR}})$ subject to the normalization $\mathbb{E}_{G_{t}^{\mathrm{FR}}}[(\alpha_{t}-\mathbb{E}_{G_{t}^{\mathrm{FR}}}[\alpha_{t}])^{2}]=1$. Intuitively, the FR flow increases probability mass at locations with high fitness but cannot create mass at new locations or move existing support points. In contrast, the W gradient flow evolves according to the continuity (transport) equation
@@
where the velocity field $V_{t}:\Theta\to\mathbb{R}^{{d_{\theta}}}$ describes the movement of mass and is chosen to maximize the gain in $F_{N}$ under the normalization $\mathbb{E}_{G_{t}^{\mathrm{W}}}[\|V_{t}\|^{2}]=1$. This trajectory effectively moves the support points across the parameter space $\Theta$.
Yan, Wang, and Rigollet (2024) show for the Gaussian location mixture case that the FR flow is guaranteed to reach a global maximum when initialized with full support, while $W$ may be stuck at a local maximum. To also exploit the gains from transporting mass across space, Yan, Wang, and Rigollet (2024) consider a discrete-time Wasserstein-Fisher-Rao (WFR) gradient flow algorithm that combines both reweighting of FR and transport dynamics of W when updating $G_{t}^{\mathrm{WFR}}$:
@@
The algorithm adaptively relocates support points so that $m$ remains moderate and need not exceed the sample size $N$. It thus inherits the global convergence guarantees of the FR flow, and adaptivity of W through mass transport. Again in a Gaussian location mixture setting, WFR is shown to be superior to the EM algorithm, and gradient-descent methods based solely on the FR or W flow. Motivated by these encouraging results, we extend the WFR from a Gaussian location mixture to the HC model.